40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 14 17:30:18 2016, ericq, Update, General, PLL measurement ongoing 
    Reply  Fri Apr 15 03:02:44 2016, ericq, Update, General, New AUX laser measurements aux_innolight_freqPSD.pdfaux_innolight_AMPM.pdfaux_innolight_data.zip
       Reply  Fri Apr 15 18:35:57 2016, gautam, Update, General, New AUX laser measurements PSL_AUX_TEMP_SCAN.pdfPSL_AUX_TEMP_SCAN.mat
          Reply  Fri Apr 15 23:11:49 2016, gautam, Update, General, Innolight 1W moved to SP table 
             Reply  Mon Apr 18 00:29:00 2016, gautam, Update, General, Beam profiling + injection current scan 6x
Message ID: 12076     Entry time: Thu Apr 14 17:30:18 2016     Reply to this: 12077
Author: ericq 
Type: Update 
Category: General 
Subject: PLL measurement ongoing 

Just a heads up that some equipment is hooked up at the PSL table for the repaired AUX laser PLL measurement, I plan to continue with it tonight.

I've taken a few spectra that, along with the PZT coefficient from the repair sheet, that suggest the noise level is ok (incoherent sum of AUX and PSL at about ~3e4 / f Hz/rtHz), but calibrated plots, etc. will follow in time.

ELOG V3.1.3-